Find the measurement of the angle of refraction using the given illustration.

N

\ |

Air (n=1.00) \ |

\ |

_____________ \ | _____________

Crown |

Glass (n= 1.52) |

|


1
Expert's answer
2022-04-24T17:19:51-0400

"\frac{\sin\alpha}{\sin\beta}=\frac{n_2}{n_1},"

"\beta=\arcsin{\frac{n_1\sin\alpha}{n_2}}=34.7°."